OpenStudy (wholock):

I need some quick help on this one question. A team of engineers must design a fuel tank in the shape of a cone. The surface area of a cone (excluding the base) is given by the formula S = pi*sqrt(r^2 + h^2) Find the radius of a cone with a height of 21 meters and a surface area of 155 meters squared. Round your answer to the nearest tenth if necessary.

OpenStudy (wholock):

@AloneS @Elsa213

OpenStudy (wholock):

@SeanStewart16

OpenStudy (seanstewart16):

osht dude, was not expecting that XD ill try, but cones beat me up

OpenStudy (wholock):

LOL, sorry, it's the reason I can here, question was beating me up.

OpenStudy (wholock):

Seriously, if you can just help me set up the equation, I could probably get it from there.

OpenStudy (johnweldon1993):

You're given everything you need, just a bit of algebra to solve for 'r' \[\large S = \pi \times \sqrt{r^2 + h^2}\] Divide both sides by \(\large \pi\) \[\large \frac{S}{\pi} = \sqrt{r^2 + h^2}\] Square both sides \[\large (\frac{S}{\pi})^2 = r^2 + h^2\] Subtract \(\large h^2\) from both sides \[\large (\frac{S}{\pi})^2 - h^2 = r^2\] And take the square root of both sides to solve for 'r' \[\large r = \sqrt{(\frac{S}{\pi})^2 - h^2}\] Now just plug in the 'S' they gave you *155* and the 'h' they gave *21* and solve for 'r

OpenStudy (wholock):

Thank you so much man! Life saver!

OpenStudy (johnweldon1993):

No problem!